The helium-neon lasers most commonly used in student physics laboratories have average power outputs of 0.250 mW.

(a) If such a laser beam is projected onto a circular spot 3.96 mm in diameter, what is its intensity?

W/m2



(b) Find the peak electric field strength.

V/m



(c) Find the peak magnetic field strength.
Answer #1

Intensity = Power / Area = 20.29 W/m2

I= c\epsilon_o E2/2 ==> E = 123.65 N/C

B=E/c = 4.12 x 10-7 Tesla

Hope you will find the answer useful. Please rate. cheers.